B4K Failure

Sunday, January 7, 2018 3:06 PM

Problem(s) Encountered:

The B4K or related electronics seems to be on the fritz.  Regardless of the type of exposure (dark, comp, etc.) the readout of the camera is three quadrants with pixel values of around 420 in ds9 and the fourth with values around 7200.  I messed with the connection between the HE and the camera to ensure it was truly connected, as this has been an issue in the past. I also power cycled the HE. My final test was to leave the HE powered down for several hours and let it cool down (even though it wasn''t particularly warm) and try again, which also failed to work.


B4K HE & cable have been sent to OSU to determine if a problem can be found.

Update 03/06/18:

After completing the usual DC voltage level checks and then half way through checking the clocks, I discovered a 7805 regulator and a 240 buffer chip on the CBB running very hot. Tracing the Ins and Outs of the buffer and looking at them on a scope showed that the Postamp control lines were being pulled down by the load (the Postamp cards). I traced the problem to postamp CCD4 which had the +15V shorted. We repaired that short (a bad bypass cap), replaced all the parts that had been running over hot and put it back together. TV mode visually looks like a working system.

Both ends of the Detector cable were damaged. The existing coax wires were move to new connector housings, backshell/clamps reassembled and beeped out to check it matched as-found condition and also Bruce''s documentation.

I blew out the Detector cable connector on the side of the HE and inspected all the coax pins which were fine.


System back in service!